# Inductive Freeness of Ziegler’s Canonical Multiderivations for Restrictions of Reflection Arrangements

### v2

@@ Abstract.
Let ${\mathscr{A}}$ be a free hyperplane arrangement. In 1989, Ziegler showed that the restriction ${\mathscr{A}}^{\prime\prime}$ of ${\mathscr{A}}$ to any hyperplane endowed with the natural multiplicity $\kappa$ is then a free multiarrangement. Recently, in [HR24], an analogue of Ziegler’s theorem for the stronger notion of inductive freeness was proved: if ${\mathscr{A}}$ is inductively free, then so is the free multiarrangement $({\mathscr{A}}^{\prime\prime},\kappa)$.
In [HR18], all reflection arrangements which admit inductively free Ziegler restrictions are classified. The aim of this paper is an extension of this classification to all restrictions of reflection arrangements utilizing the aforementioned fundamental result from [HR24].
@@ 1. Introduction
The class of free arrangements, respectively free multiarrangements, plays a fundamental role in the theory of hyperplane arrangements, respectively multiarrangements. In [Z89], Ziegler introduced the notion of multiarrangements and initiated the study of their freeness. We begin by recalling Ziegler’s fundamental construction from loc. cit.
@@
Because of the relevance of Ziegler’s multiplicity in the theory of free arrangements, it is natural to investigate stronger freeness properties for $({\mathscr{A}}^{\prime\prime},\kappa)$ and specifically to ask for an analogue of Theorem 1.2 for inductive freeness. This was settled recently in [HR24]:
Theorem 1.3 ([HR24, Thm. 1.3]).
@@
Even if ${\mathscr{A}}$ itself is not inductively free, the following result from [HR24] still guarantees inductive freeness of $({\mathscr{A}}^{\prime\prime},\kappa)$, provided the restriction ${\mathscr{A}}^{\prime\prime}$ itself is inductively free.
Theorem 1.4 ([HR24, Thm. 1.6]).
@@
We note that here the freeness assumption on ${\mathscr{A}}\setminus\{H_{0}\}$ is needed, cf. [HR24, Ex. 4.2].
@@
In [HR18], all reflection arrangements ${\mathscr{A}}(W)$ were classified which admit Ziegler restrictions $({\mathscr{A}}(W)^{\prime\prime},\kappa)$ that are inductively free.
@@
The reverse implication in part (i) of Theorem 1.7 follows from Theorem 1.3 and the fact that ${\mathscr{A}}(W)$ is inductively free if and only if it is hereditarily inductively free, thanks to [HR15, Thm. 1.2]; see Theorem 3.1(ii).
@@
We mention that the equivalences from both Corollaries 1.6 and 1.8 are false in general. For the failure of the forward implication, see [HR24, Ex. 1.13] and for the failure of the reverse implication, see [HR18, Ex. 2.24]. In view of these elementary counterexamples, the equivalences of these two corollaries are rather striking.
We also study multiplicities which are concentrated at a single hyperplane. These were introduced by Abe, Terao and Wakefield, [ATW08, §5]. It turns out that they are closely related to Ziegler’s canonical multiplicity, see Proposition 5.1.
@@
The forward implication of Theorem 1.11 is [HR24, Cor. 5.6]. We observe that Theorem 1.11 gives a uniform case free proof of [HR18, Thm. 1.9].
@@ 2.2. Freeness of multiarrangements
We also require the following tool from [HR24]. This result turns out to be useful in our sequel in determining inductively free Ziegler restrictions.
Lemma 2.8 ([HR24, Lem. 2.12]).
@@
We recall a natural partial order on the set of multiplicities for a simple arrangement and the notion of an additively free multiarrangement from [HR24, Def. 2.19].
@@ 3. Reflection arrangements and their restrictions
Theorem 1.7 and Corollary 1.8 follow from all of the subsequent lemmas along with Theorem 3.4 and Proposition 3.3.
@@ 4.1. The monomial groups $G(r,r,n)$
It follows from Proposition 3.3(ii) that if we restrict ${\mathscr{A}}^{k}_{\ell}(r)$ to a coordinate hyperplane, then we obtain ${\mathscr{A}}^{\ell-1}_{\ell-1}(r)={\mathscr{A}}(G(r,1,\ell-1))$. We consider that case first, cf. [HR24, Ex. 1.14].
@@
Next we consider the instances when ${\mathscr{A}}={\mathscr{A}}_{\ell}^{k}(r)$ is not inductively free, i.e., when $31\leq k\leq\ell-$ 3, cf. Theorem 3.4(i)(b) (cf. [HR24, Ex. 1.14]).
@@
Table 2. Induction table for ${\mathscr{A}}={\mathscr{A}}^{l-4}_{l}(r),({\mathscr{A}}^{H},\kappa)=({\mathscr{A}}^{l-3}_{l-1}(r),\kappa)$
@@
Owing to Corollary 4.1 and Lemma 4.3, for ${\mathscr{A}}={\mathscr{A}}_{\ell}^{k}(r)$ with $r\geq 3$, $\ell\geq 5$ and $31\leq k\leq\ell-$ 3, it depends on the choice of $H_{0}$ whether $({\mathscr{A}}^{\prime\prime},\kappa)$ is inductively free or not.
@@ 4.2. Groups of type $G_{33}$ and $G_{34}$
To not interrupt the present flow, we defer the proof of the lemma to an appendix.
@@
Recall the notion of a concentrated multiplicity $\delta$ from Definition 1.9. The following combines [ATW08, Prop. 5.2], parts of its proof and Theorem 1.2. While not needed for the proof of Theorem 1.11, this result does illustrate the connection with the canonical multiplicity.
Proposition 5.1 ([HR18, Prop. 2.14]).
  • (i)
  • (ii)
@@
Acknowledgments: This work was supported by DFG-Grant RO 1072/21-1 (DFG Project number 494889912) to G. Röhrle.

### v3

@@ Abstract.
Let ${\mathscr{A}}$ be a free hyperplane arrangement. In 1989, Ziegler showed that the restriction ${\mathscr{A}}^{\prime\prime}$ of ${\mathscr{A}}$ to any hyperplane endowed with the natural multiplicity $\kappa$ is then a free multiarrangement. Recently, in [HR25], an analogue of Ziegler’s theorem for the stronger notion of inductive freeness was proved: if ${\mathscr{A}}$ is inductively free, then so is the free multiarrangement $({\mathscr{A}}^{\prime\prime},\kappa)$.
In [HR18], all reflection arrangements which admit inductively free Ziegler restrictions are classified. The aim of this paper is to extend this classification to all arrangements which are induced by reflection arrangements utilizing the aforementioned fundamental result from [HR25].
@@ 1. Introduction
1.1.
The class of free arrangements, respectively free multiarrangements, plays a fundamental role in the theory of hyperplane arrangements, respectively multiarrangements. In [Z89], Ziegler introduced the notion of multiarrangements and initiated the study of their freeness. We begin by recalling Ziegler’s fundamental construction from [Z89].
Let ${\mathbb{K}}$ be a field and let $V$ be a finite-dimensional ${\mathbb{K}}$-vector space. A hyperplane arrangement or simply an arrangement is a pair $({\mathscr{A}},V)$, where ${\mathscr{A}}$ is a finite collection of (central) hyperplanes in $V$. Usually, we just write ${\mathscr{A}}$ in place of $({\mathscr{A}},V)$; see Section 2.1.
@@
Because of the relevance of Ziegler’s multiplicity in the theory of free arrangements, it is natural to investigate stronger freeness properties for $({\mathscr{A}}^{\prime\prime},\kappa)$ and specifically to ask for an analogue of Theorem 1.2 for inductive freeness. This was settled recently in [HR25]:
Theorem 1.3 ([HR25, Thm. 1.3]).
@@
Even if ${\mathscr{A}}$ itself is not inductively free, the following result from [HR25] still guarantees inductive freeness of $({\mathscr{A}}^{\prime\prime},\kappa)$, provided the restriction ${\mathscr{A}}^{\prime\prime}$ itself is inductively free.
Theorem 1.4 ([HR25, Thm. 1.6]).
@@
We note that here the freeness assumption on ${\mathscr{A}}\setminus\{H_{0}\}$ is needed, cf. [HR25, Ex. 4.2].
@@
1.2.
In this subsection assume that ${\mathbb{K}}={\mathbb{C}}$, the complex numbers, and let $W$ be an irreducible complex reflection group with $V$ the ${\mathbb{C}}$-vector space affording the reflection representation of $W$. Denote by ${\mathscr{A}}(W)$ the reflection arrangement of $W$ in $V$ which consists of all the complex reflecting hyperplanes of $W$. In [HR18], all reflection arrangements ${\mathscr{A}}(W)$ were classified which admit Ziegler restrictions $({\mathscr{A}}(W)^{\prime\prime},\kappa)$ that are inductively free:
@@
We mention that the equivalences from both Corollaries 1.6 and 1.8 are false in general (i.e., if ${\mathscr{A}}$ is not a reflection arrangement, and not a restriction of such). For the failure of the forward implication, see [HR25, Ex. 1.13] and for the failure of the reverse implication, see [HR18, Ex. 2.24]. In view of these elementary counterexamples, the equivalences of these two corollaries are rather striking.
1.3.
In this subsection we allow ${\mathbb{K}}$ to be an arbitrary field again. In a separate part, we also study multiplicities which are concentrated at a single hyperplane. These were introduced by Abe, Terao and Wakefield, [ATW08, §5]. It turns out that they are closely related to Ziegler’s canonical multiplicity, see Proposition 1.12.
@@
The forward implication of Theorem 1.11 is [HR25, Cor. 5.6]. We observe that Theorem 1.11 gives a uniform case free proof of [HR18, Thm. 1.9].
The following combines [ATW08, Prop. 5.2], parts of its proof and Theorem 1.2. While not needed for the proof of Theorem 1.11, this result does illustrate the connection with the canonical multiplicity.
Proposition 1.12 ([HR18, Prop. 2.14]).
  • (i)
  • (ii)
@@ 2.2. Freeness of multiarrangements
We also require the following tool from [HR25]. This result turns out to be useful in our sequel in determining inductively free Ziegler restrictions.
Lemma 2.8 ([HR25, Lem. 2.12]).
@@
We recall a natural partial order on the set of multiplicities for a simple arrangement and the notion of an additively free multiarrangement from [HR25, Def. 2.19].
@@ 4.1. The monomial groups $G(r,r,n)$
It follows from Proposition 3.3(ii) that if we restrict ${\mathscr{A}}^{k}_{\ell}(r)$ to a coordinate hyperplane, then we obtain ${\mathscr{A}}^{\ell-1}_{\ell-1}(r)={\mathscr{A}}(G(r,1,\ell-1))$. We consider that case first, cf. [HR25, Ex. 1.14].
@@
Next we consider the instances when ${\mathscr{A}}={\mathscr{A}}_{\ell}^{k}(r)$ is not inductively free, i.e., when $31\leq k\leq\ell-$ 3, cf. Theorem 3.4(i)(b) (cf. [HR25, Ex. 1.14]).
@@
Table 2. Induction table for ${\mathscr{A}}={\mathscr{A}}^{\ell-4}_{\ell}(r),({\mathscr{A}}^{H},\kappa)=({\mathscr{A}}^{\ell-3}_{\ell-1}(r),\kappa)$
@@
We conclude this subsection with the following observation.
@@
Owing to the results above, for ${\mathscr{A}}={\mathscr{A}}_{\ell}^{k}(r)$ with $r\geq 3$, $\ell\geq 5$ and $31\leq k\leq\ell-$ 3, it depends on the choice of $H_{0}$ whether $({\mathscr{A}}^{\prime\prime},\kappa)$ is inductively free or not (e.g., see Corollary 4.1 and Lemma 4.3).
@@ 4.2. Groups of type $G_{33}$ and $G_{34}$
To not interrupt the present flow, we defer the proof of the lemma to the appendix.
4.3. Proof for Theorem 1.7
We finally present the proof of our main theorem. Thanks to Theorems 1.3 and 3.4, for the reverse implication of Theorem 1.7 we only need to investigate the instances when either ${\mathscr{A}}(G(r,r,n))^{X}\cong{\mathscr{A}}^{k}_{\ell}(r)$, where $\ell=\dim X$ and $31\leq k\leq\ell-$ 3, or when $W$ is of type $G_{33}$, or $G_{34}$ and $X\in L({\mathscr{A}}(W))\setminus\{V\}$ with $\dim X\geq 4$.
The reverse implication in part (i) of Theorem 1.7 follows from Theorem 3.2.
The reverse implication in part (ii) of Theorem 1.7 follows from Corollary 4.1, and Lemmas 4.2, 4.6, and 4.7 along with Theorem 3.4(b) and Proposition 3.3(b). While the forward implication in part (ii) follows from Lemma 4.3 together with Theorem 3.4(b) and Proposition 3.3(b).
For part (iii) of Theorem 1.7, the equivalence follows from Lemmas 4.9 and 4.10.
For the forward implication of Theorem 1.7, assume that we are not in case (i). Then by Theorem 3.4, we are in case (ii) or case (iii). So we are done.
@@
Acknowledgments: This work was supported by DFG-Grant RO 1072/21-1 (DFG Project number 494889912) to G. Röhrle. We thank the anonymous referees for several suggestions which have improved the exposition.